# Hamiltonian stochastic Lie systems and applications

### v2

@@ 7 Stochastic Poisson coalgebra method
Let us review the Poisson coalgebra method for Hamiltonian systems relative to different geometric structures. In general, our approach slightly matches what can be found in previous works, but it also introduces certain simplifications and assumptions that allow one to apply it to many topics. In particular, the method can still be applied to Hamiltonian stochastic Lie systems by considering that they are determined by a certain family of Hamiltonian functions. Indeed, our present procedure is a new modification of the coalgebra method for deriving superposition rules for $k$-symplectic Lie systems devised in [24].
@@
    Consider a Vessiot–Guldberg Lie algebra spanned by a basis of vector fields $X_{1},\ldots,X_{r}$ on the manifold $M$ where the stochastic Lie system is defined on.
@@
    Use local coordinates $x^{1},\ldots,x^{n}$ on $M$ and consider this coordinate system to be defined on each copy of $M$ within $M^{m+1}$ to get a coordinate system $\{x^{i}_{(a)}\mid i=1,\ldots,n,\,\,a=0,\ldots,m\}$ on $M^{m+1}$. Obtain $n$ functionally independent first-integrals $F_{1},\ldots,F_{n}$ common to all diagonal prolongations $X^{[m+1]}_{1},\ldots,X^{[m+1]}_{r}$ such that
@@
    The condition (7.1) allows us to ensure that the equations $F_{i}=k_{i}$, for $i=1,\ldots,n$, enable us to write the expressions of the variables $x_{(0)}^{1}\ldots,x_{(0)}^{n}$ in terms of $x^{1}_{(a)},\ldots,x_{(a)}^{n}$, with $a=1,\ldots,m,$ and $k_{1},\ldots,k_{n}$.
@@
Note that every Hamiltonian stochastic Lie system is related to a Stratonovich operator $\mathfrak{H}$ given by $\ell$ components and each one can be understood as a $t$-dependent vector field (that can be also $\mathbb{R}^{\ell}$-dependent if we allow the dependence on all stochastic variables). Hence, $\mathfrak{H}$ can be understood as an $t$-dependent $\ell$-vector field, i.e. for every $t\in\mathbb{R}$, it is a section of the $\ell$-tangent bundle of velocities $\mathrm{T}^{\ell}M=\mathrm{T}M\oplus\dotsb\oplus\mathrm{T}M\rightarrow M$, where $\mathrm{T}M\oplus\dotsb\oplus\mathrm{T}M$ is to be considered as the Whitney sum of the tangent bundle to $M$ with itself. In this manner, one can prolong it to a section $\mathfrak{H}$ of ×$\mathbb{R}\times(\mathrm{T}^{k}M)^{[m]}\rightarrow\mathbb{R}\times M^{m}$. Moreover, $\mathfrak{H}$ is related to a certain $t$-dependent Hamiltonian function $h\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$. Then, $h$ is a section of ×$\mathbb{R}\times M\times\mathbb{R}^{\ell}\rightarrow\mathbb{R}\times M$ and it can be extended to a section of ×$\mathbb{R}\times M^{m}\times\mathbb{R}^{\ell}\rightarrow\mathbb{R}\times M^{m}$. Moreover, the prolongation $h^{[m]}$ is the $t$-dependent Hamiltonian function related to $\mathfrak{H}^{[m]}$. Then, the following result is immediate.
@@
If $\mathfrak{H}$ is a Hamiltonian stochastic Lie system admitting a $t$-dependent Hamiltonian $h\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ relative to a symplectic form $\omega$, then $\mathfrak{H}^{[m]}$ is a Hamiltonian stochastic Lie system relative to the symplectic form $\omega^{[m]}$ admitting a Hamiltonian $h^{[m]}\colon\mathbb{R}\times M^{m}\rightarrow\mathbb{R}^{\ell}$. In particular, if $h_{1},\ldots,h_{r}$ is a basis of the Lie algebra of Hamilton functions for $\mathfrak{H}$, then $h_{1}^{[m]},\ldots,h_{r}^{[m]}$ is a basis of the Lie algebra of Hamiltonian functions for $\mathfrak{H}^{[m]}$.
One has the following immediate result.
@@
The space of $\ell$-Hamiltonian functions on a symplectic manifold $(M,\omega)$, with Poisson bracket {,}$\{\cdot,\cdot\}_{\omega}$ induced by $\omega$, is a Poisson algebra relative to the bracket
$\{h,h^{\prime}\}_{\ell}=(\{h_{1},h^{\prime}_{1}\}_{\omega},\ldots,\{h_{\ell},h^{\prime}_{\ell}\}_{\omega})\,,\qquad\forall h=(h_{1},\ldots,h_{\ell}),h^{\prime}=(h^{\prime}_{1},\ldots,h^{\prime}_{\ell})\in\mathscr{C}^{\infty}(M,\mathbb{R}^{\ell})\,,$
@@
It is worth recalling that the space of $\ell$-Hamiltonian functions relative to an $\ell$-symplectic structure was not a Poisson algebra due to the fact that the multiplication of $\ell$-Hamiltonian functions could not be ensured to be an $\ell$-Hamiltonian function as each coordinate is assumed to be a Hamiltonian function of the same vector field and this cannot be ensured for their bracket as each coordinate of the bracket is associated with different presymplectic forms and may be related to different vector fields [24]. Here, this problem does not appear, since each $h_{i}$, with $i=1,\ldots,\ell$, may be the Hamiltonian function of a different vector field.
For the sake of completeness, let us prove the following result.
@@
Let $\mathfrak{H}$ be a Hamiltonian stochastic Lie system with respect to a symplectic form $\omega$ and possessing a smallest (in the sense of inclusion) Lie–Hamiltonian Vessiot–Guldberg Lie algebra $(\mathfrak{W},\{h_{1},\ldots,h_{r}\})$ relative to the Poisson bracket related to $\omega$. A function $f\in\mathscr{C}^{\infty}(M)$ is a strong constant of the motion for $\mathfrak{H}$ if commutes with all the elements of ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}},\{\cdot,\cdot\})$. In particular, $f$ is a strong contact of the motion if it commutes with the elements of $\mathfrak{M}$.
@@
The function $f$ is a strong constant of the motion for $\mathfrak{H}$ if
@@
where the $X_{t}^{i}$ are the components of the $t$-dependent $\ell$-vector field related to $\mathfrak{H}$. Since $X_{t}^{i}=\{h^{i}_{t},f\}$ for $i=1,\ldots,\ell$, the result follows. Note also that
@@
Inductively, $f$ is shown to commute with all the elements of ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}})$. In particular, ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}})\subset\mathfrak{W}$ ensures that $f$ is a strong constant of motion.
In Lie–Hamilton systems, a $t$-dependent Hamiltonian vector field admitted for every $t\in\mathbb{R}$ a Hamiltonian function belonging to a finite-dimensional Lie algebra of Hamiltonian functions relative to a Poisson bracket.
@@
Let $\mathfrak{H}$ be a Hamiltonian stochastic Lie system with an associated Lie–Hamilton Vessiot–Guldberg Lie algebra $(\mathfrak{W}=\langle h_{1},\ldots,h_{r}\rangle,\{\cdot,\cdot\}_{\omega})$ relative to the symplectic form $\omega$. Let $\{v_{1},\ldots,v_{r}\}$ be a basis of linear coordinates on $\mathfrak{g}^{*}\simeq\mathfrak{W}$. Given the momentum map $J:M\to\mathfrak{g}^{\ast}$, the pull-back $J^{\ast}C$ of any Casimir function $C$ on $\mathfrak{g}^{\ast}$ is a constant of the motion for $\mathfrak{H}$. Moreover, if $h_{i}=J^{*}v^{i}$ for $i=1,\ldots,r$, and $C=C(v_{1},\ldots,v_{r})$, then
$C\left(\sum_{a=1}^{m}h_{1}(x_{(a)}),\ldots,\sum_{a=1}^{m}h_{r}(x_{(a)})\right)\,,$ (7.3)
is a constant of the motion of $\mathfrak{H}^{[m]}$.
The stochastic Poisson coalgebra method takes its name from the fact that it is applied to Hamiltonian stochastic Lie systems and it analyses the use of Poisson coalgebras and a so-called coproduct to obtain superposition rules. In fact, the coproduct is responsible for the form of (7.3) (see [7]).
@@ 8 Conclusions and outlook
This paper has reviewed in a simple manner the theory of stochastic Lie systems solving a mistake in the literature. Some relations between the Stratonovich and the Itô approaches to stochastic Lie systems have been reviewed, which has shown that the stochastic Lie theorem in the Itô approach has a different non-expected form. Types of Hamiltonian stochastic Lie systems have been introduced as well as some possible generalisations. We have also introduced the theory of Hamilton stochastic Lie systems. The theory of stability and relative equilibrium points for Hamiltonian stochastic differential equations has been developed and, in particular, we have focused in the study of Hamiltonian stochastic Lie systems relative to symplectic forms. Some results concerning the energy-momentum method for stochastic Lie systems have been obtained, generalizing [5]. Many examples concerning physical/epidemiological applications have been developed. We have also remarked that the theory of Poisson coalgebras can be developed for our Hamiltonian stochastic Lie systems.
@@
Concerning applications, stochastic SIS models have been analysed. Such models use to be studied using deterministic methods in the literature. Instead, we here approach them in a pure stochastic approach. Many other physical systems have also been analysed.
@@
In the future, we aim to further develop the energy-momentum method for stochastic Lie systems. This entails the search for criteria ensuring our of a Hamiltonian Stratonovich operator the stability of equilibrium points obtained after reduction. Moreover, we expect, as in the deterministic case, to obtain certain degeneracies that will have to be analysed to solve the problem. Note that a further study of the advantages of epidemiological models from a pure stochastic point of view will have to be addressed. Moreover, we also plan to study a stochastic Hamiltonian system appearing in celestial mechanics concerning the stochastic variation of the inertia tensor [52]. Finally, it would be interesting to study other types of superposition rules appearing in stochastic differential equations [37] and applying the Poisson coalgebra method for Hamiltonian stochastic Lie systems .

### v3

@@ 7 Stochastic Poisson coalgebra method
Let us review and extend the Poisson coalgebra method for Hamiltonian systems relative to a symplectic form. In general, our theory is a stochastic generalisation of what can be found in the classical setting [26, Section 4.2.7]. Although the procedure is very similar to the original approach, a few key differences allow its application to many new domains. In particular, the method can still be applied to Hamiltonian stochastic Lie systems by considering that they are determined by a certain $\ell$-family of $t$-dependent Hamiltonian functions. Indeed, our procedure is a new modification of the coalgebra method for deriving superposition rules for $k$-symplectic Lie systems (see [26, Section 7.8] and references therein).
@@
    Consider a Vessiot–Guldberg Lie algebra of the stochastic Lie system spanned by a basis of vector fields $X_{1},\ldots,X_{r}$ on the manifold $M$.
@@
    Use local coordinates $x^{1},\ldots,x^{n}$ on $M$ and consider this coordinate system to be defined on each copy of $M$ within $M^{m+1}$ to get a coordinate system $\{x^{i}_{(a)}\mid i=1,\ldots,n,\,\,a=0,\ldots,m\}$ on $M^{m+1}$. Obtain first integrals $F_{1},\ldots,F_{n}$ common to all the diagonal prolongations $X^{[m+1]}_{1},\ldots,X^{[m+1]}_{r}$ such that
@@
    Condition (7.1) allows us to ensure that the equations $F_{i}=k_{i}$, for $i=1,\ldots,n$, enable us to write the expressions of the variables $x_{(0)}^{1}\ldots,x_{(0)}^{n}$ in terms of $x^{1}_{(a)},\ldots,x_{(a)}^{n}$, with $a=1,\ldots,m,$ and $k_{1},\ldots,k_{n}$.
@@
Note that every Hamiltonian stochastic Lie system is related to a Stratonovich operator $\mathfrak{H}$ given by $\ell$ components and each one can be understood as a $t$-dependent vector field. Hence, $\mathfrak{H}$ can be understood as a $t$-dependent $\ell$-vector field, i.e. for every $t\in\mathbb{R}$, it is a section of the $\ell$-tangent bundle of velocities $\mathrm{T}^{\ell}M=\mathrm{T}M\oplus\dotsb\oplus\mathrm{T}M\rightarrow M$, where $\mathrm{T}M\oplus\dotsb\oplus\mathrm{T}M$ is to be considered as the Whitney sum of the tangent bundle to $M$ with itself. Similarly, one can prolong diagonally $\mathfrak{H}$ to a map $\mathfrak{H}^{[m]}:\mathrm{T}\mathbb{R}^{\ell}\times M^{m}\rightarrow\mathrm{T}M^{m}$. Moreover, $\mathfrak{H}$ is related to a $t$-dependent Hamiltonian function $h\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ which gives rise, by prolonging diagonally each of its components for every fixed $t$, to a mapping $h^{[m]}:\mathbb{R}\times M^{m}\rightarrow\mathbb{R}^{\ell}$. Additionally, the diagonal prolongation $h^{[m]}$ is the $t$-dependent Hamiltonian function related to $\mathfrak{H}^{[m]}$. Then, the following result is immediate.
@@
If $\mathfrak{H}$ is a Hamiltonian stochastic Lie system admitting a $t$-dependent Hamiltonian $h\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ relative to a symplectic form $\omega$, then $\mathfrak{H}^{[m]}$ is a Hamiltonian stochastic Lie system relative to the symplectic form $\omega^{[m]}$ admitting a $t$-dependent Hamiltonian $h^{[m]}\colon\mathbb{R}\times M^{m}\rightarrow\mathbb{R}^{\ell}$. In particular, if $h_{1},\ldots,h_{r}$ is a basis of a Lie–Hamilton Lie algebra for $\mathfrak{H}$, then $h_{1}^{[m]},\ldots,h_{r}^{[m]}$ is a basis of a Lie–Hamilton Lie algebra for $\mathfrak{H}^{[m]}$.
We have the following immediate proposition.
@@
The space of $\ell$-Hamiltonian functions, namely $\mathscr{C}^{\infty}(M,\mathbb{R}^{\ell})$ on a symplectic manifold $(M,\omega)$, with Poisson bracket $\{\cdot,\cdot\}_{\omega}$ induced by $\omega$, is a Poisson algebra relative to the bracket
$\{h,h^{\prime}\}_{\ell}=(\{h_{1},h^{\prime}_{1}\}_{\omega},\ldots,\{h_{\ell},h^{\prime}_{\ell}\}_{\omega})\,,\quad\forall h=(h_{1},\ldots,h_{\ell}),\quad h^{\prime}=(h^{\prime}_{1},\ldots,h^{\prime}_{\ell})\in\mathscr{C}^{\infty}(M,\mathbb{R}^{\ell})\,$,
@@
It is worth recalling that the space of $\ell$-Hamiltonian functions relative to an $\ell$-symplectic form was not a Poisson algebra. This was due to the fact that the multiplication of $\ell$-Hamiltonian functions could not be ensured to be an $\ell$-Hamiltonian function as each coordinate is assumed to be a Hamiltonian function of the same vector field and this cannot be ensured for their bracket as each coordinate of the bracket is associated with different presymplectic forms and may be related to different vector fields (see [26, Chapter 7]). Here, this problem does not appear, since each $h_{i}$, with $i=1,\ldots,\ell$, may be the Hamiltonian function of a different vector field.
For the sake of completeness, let us prove the following result. Recall that ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}},\{\cdot,\cdot\}_{\omega})$ is the smallest Lie algebra of Hamiltonian functions relative to $\omega$ containing $\{h_{t}\}_{t\in\mathbb{R}}$.
@@
Let $\mathfrak{H}$ be a Hamiltonian stochastic Lie system with respect to a symplectic form $\omega$ and possessing a Lie–Hamilton Lie algebra $\mathfrak{W}=\langle h_{1},\ldots,h_{r}\rangle$ relative to the Poisson bracket related to $\omega$. A function $f\in\mathscr{C}^{\infty}(M)$ is a strong constant of motion for $\mathfrak{H}$ if commutes with all the elements of ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}},\{\cdot,\cdot\}_{\omega})$. In particular, $f$ is a strong contact of the motion if it commutes with the elements of $\mathfrak{M}$.
@@
The function $f$ is a strong constant of motion for $\mathfrak{H}$ if
@@
where the $X_{t}^{i}$ are the components at a fixed $t$ of the $t$-dependent $\ell$-vector field related to $\mathfrak{H}$. Since $X_{t}^{i}f=\{h^{i}_{t},f\}_{\omega}$ for $i=1,\ldots,\ell$ and $t\in\mathbb{R}$, the result follows. Note also that
@@
Inductively, $f$ is commutes with all the elements of ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}},\{\cdot,\cdot\}_{\omega})$. Since we restrict ourselves to the case for a symplectic form $\omega$, one has that ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}},\{\cdot,\cdot\}_{\omega})$ is included in $\mathfrak{W}\oplus\mathbb{R}$, where $\mathbb{R}$ stands for the space of constant functions. And the latter ensures that $f$ is a strong constant of motion.
In Lie–Hamilton systems, a $t$-dependent Hamiltonian vector field admits for every $t\in\mathbb{R}$ a Hamiltonian function belonging to a finite-dimensional Lie algebra of Hamiltonian functions relative to a Poisson bracket.
@@
Let $\mathfrak{H}$ be a Hamiltonian stochastic Lie system with an associated Lie–Hamilton Lie algebra $(\mathfrak{W}=\langle h_{1},\ldots,h_{r}\rangle,\{\cdot,\cdot\}_{\omega})$ relative to the symplectic form $\omega$. Let $\{v_{1},\ldots,v_{r}\}$ be a basis of linear coordinates on $\mathfrak{g}^{*}\simeq\mathfrak{W}$. Given the momentum map $J:\Gamma\in M\mapsto h_{i}(\Gamma)=(J^{*}v^{i})(\Gamma)\in\mathfrak{g}^{\ast}$, the pull-back $J^{\ast}C$ of any Casimir function $C$ on $\mathfrak{g}^{\ast}$ is a constant of motion for $\mathfrak{H}$. Moreover, if $C=C(v_{1},\ldots,v_{r})$, then
$C\left(\sum_{a=1}^{k}h_{1}(x_{(a)}),\ldots,\sum_{a=1}^{k}h_{r}(x_{(a)})\right)\,,\qquad 1\leq k\leq m,$ (7.3)
is a constant of motion of $\mathfrak{H}^{[m]}$.
The stochastic Poisson coalgebra method above takes its name from the fact that it is applied to Hamiltonian stochastic Lie systems and analyses the use of Poisson coalgebras and a so-called coproduct to obtain superposition rules. In fact, the coproduct is responsible for the form of (7.3) (see [26]). Although we have focused on Hamiltonian symplectic systems, the above Poisson coalgebra method can also be applied to Hamiltonian systems relative to many other geometric structures.
Let us provide a simple example illustrating the above techniques based on the system
$\delta\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}=\left[\begin{pmatrix}0&1\\ -\omega^{2}(t)&0\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}+\begin{pmatrix}0\\ g(t)\end{pmatrix}\right]\delta t+\begin{pmatrix}0\\ \alpha(t)\end{pmatrix}\circ\delta W_{1},$ (7.4)
for a Brownian motion $W_{1}$, and any $t$-dependent functions $g(t)$, $\omega(t)$, and $\alpha(t)$, which retrieves certain oscillators in [25] and is connected to numerous physical oscillators with driven stochastic components. Since
$\mathfrak{H}=(\omega^{2}(t)X_{3}+X_{1}+g(t)X_{5},\alpha(t)X_{5}),$
system (7.4) admits a Vessiot–Guldberg Lie algebra of Hamiltonian vector fields $V_{o}$ relative to $\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$ spanned by
$X_{1}=\sum_{i=1}^{n}y_{i}\frac{\partial}{\partial x_{i}}\,,\qquad X_{2}=\frac{1}{2}\sum_{i=1}^{n}\left(x_{i}\frac{\partial}{\partial x_{i}}-y_{i}\frac{\partial}{\partial y_{i}}\right)\,,\qquad X_{3}=-\sum_{i=1}^{n}x_{i}\frac{\partial}{\partial y_{i}},\qquad$
$X_{4}=\sum_{i=1}^{n}\frac{\partial}{\partial x_{i}},\qquad X_{5}=\sum_{i=1}^{n}\frac{\partial}{\partial y_{i}}.$
Their non-vanishing commutation relations read
$X_{1},X_{2}]=X_{1},\qquad[X_{1},X_{3}]=2X_{2},\qquad[X_{2},X_{3}]=X_{3},$
$X_{1},X_{5}]=-X_{4},\qquad[X_{3},X_{4}]=X_{5},\qquad[X_{2},X_{4}]=-\frac{1}{2}X_{4},\qquad[X_{2},X_{5}]=\frac{1}{2}X_{5}.$
Therefore, the Lie algebra $V_{o}$ is isomorphic to the semi-direct sum of Lie algebras $\mathfrak{sl}_{2}\ltimes\mathbb{R}^{2}$. In particular, $\langle X_{1},X_{2},X_{3}\rangle$ is isomorphic to $\mathfrak{sl}_{2}$, while $\langle X_{4},X_{5}\rangle$ is an Abelian ideal of $V_{o}$ . The vector fields $X_{1},\ldots,X_{5}$ are related to a Lie algebra of Hamiltonian functions spanned by
$h_{1}=\frac{1}{2}\sum_{i=1}^{n}y_{i}^{2},\quad h_{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}y_{i},\quad h_{3}=\frac{1}{2}\sum_{i=1}^{n}x_{i}^{2},\quad h_{4}=\sum_{i=1}^{n}y_{i},\quad h_{5}=-\sum_{i=1}^{n}x_{i},\quad h_{6}=n,$ (7.5)
Hence, they span a Lie algebra $\mathfrak{W}_{0}$ isomorphic to the so-called two-photon algebra [9]. Indeed, $\langle h_{1},h_{2},h_{3}\rangle$ is isomorphic to $\mathfrak{sl}_{2}$, while $\langle h_{4},h_{5},h_{6}\rangle$ is a three-dimensional Heisenberg algebra $\mathfrak{h}_{3}$ and the LH algebra $\mathfrak{W}_{0}$ is isomorphic to $\mathfrak{g}_{0}\simeq\mathfrak{sl}_{2}\ltimes\mathfrak{h}_{3}$. This gives rise to a momentum map $J_{n}:\Gamma\in\mathrm{T}\mathbb{R}^{n}\mapsto h_{1}(\Gamma)e^{1}+\ldots+h_{6}(\Gamma)e^{6}\in\mathfrak{g}_{0}^{*}$. Note that $\{e_{1},\ldots,e_{6}\}$ is a basis of $\mathfrak{g}_{0}$ spanning the same commutation relations than (7.5). Then, one has the Casimir of $\mathfrak{g}_{0}$ given by  [^2] The isomorphism $(h_{0},h_{1},h_{2},h_{3},h_{4},h_{5})\leftrightarrow(e_{6},-e_{5},e_{4},-2e_{2},e_{3},-e_{1})$ from the Hamiltonian functions in [12, p. 27] maps the Casimir in [12] into $C$.
$C=2(-e_{5}^{2}e_{1}-e_{4}^{2}e_{3}-2e_{5}e_{4}e_{2})-4e_{6}(e_{2}^{2}-e_{3}e_{1}),$
which gives rise to a constant of motion of (7.4) of the form
$F_{n}=J_{n}^{*}C=2(-h_{5}^{2}h_{1}-h_{4}^{2}h_{3}-2h_{5}h_{4}h_{2})-4h_{6}(h_{2}^{2}-h_{3}h_{1}).$
One can also see that $h_{1},\ldots,h_{6}$ are the diagonal prolongations to $(\mathrm{T}\mathbb{R})^{n}$ of the oscillator (7.4) for $n=1$. Then, $F_{n}$ is nothing but the constant of motion (7.3) given for this specific case. Moreover,
$F_{1}=F_{2}=0,\qquad F_{3}=(x_{3}(y_{2}-y)+x_{2}(y_{1}-y_{3})+x_{1}(y_{3}-y_{2}))^{2}$
while further functions $F_{n}$ for $n>3$ have a complicated expression.
@@ 8 Conclusions and outlook
Our paper provided an introduction to stochastic Lie systems that could draw the attention of people working on stochastic systems or Lie systems. In this respect, some basic notions and results in stochastic systems and Lie systems have been explained in detail. In particular, some types of semi-martingales, which are suitable for use in the theory, have been applied. Meanwhile, some differential geometry concepts and other ideas for Lie systems have been explained through initial examples. This paper has skipped many technical details of the introduction to stochastic Lie systems in [48] that can be omitted in practical applications. This has made our presentation more accessible to a general public. We also reviewed the theory of stochastic Lie systems, correcting a mistake in the literature to give a solid foundation for the theory. Some relations between the Stratonovich and the Itô approaches to stochastic Lie systems have been reviewed. This showed that the stochastic Lie theorem in the Itô framework is quite different from its classical form. Moreover, we proved that this fact is relevant when considering practical applications of the theory.
Types and generalisations of stochastic Lie systems have been introduced. In particular, we have focused on the study of Hamiltonian stochastic Lie systems. The theory of stability and relative equilibrium points for Hamiltonian stochastic differential equations has been developed and, in particular, we have focused in the study of Hamiltonian stochastic Lie systems relative to symplectic forms. Several examples with potential applications have been provided so as to illustrate the theory.
Some results concerning the energy-momentum method for stochastic Lie systems have been obtained, generalising [8]. Many examples concerning physical/epidemiological applications have been developed. We have also remarked that the theory of Poisson coalgebras can be developed for our Hamiltonian stochastic Lie systems.
@@
Concerning applications, stochastic SIS models have been analysed. Such models are usually studied using deterministic methods in the literature. Instead, we approach them here in a pure stochastic way. Many other physical systems, like different types of stochastic oscillators with damped terms, have also been analysed.
@@
In the future, we aim to further develop the energy-momentum method for stochastic Lie systems. This entails the search for criteria ensuring the stability of equilibrium points, for a Hamiltonian Stratonovich operator, after reduction. Moreover, we expect, as in the deterministic case, to obtain certain degeneracies of the Hamiltonian functions that will need to be analysed to solve the problem [54]. Note that a further study of the advantages of epidemiological models from a purely stochastic point of view will need to be undertaken. In particular, we plan to study SIS models that have been traditionally studied in a deterministic Hamiltonian manner [16, 29] via stochastic Lie systems, and analyse their possible advantages. There seems to be a quite large new field of applications for Hamiltonian stochastic Lie systems or Hamiltonian foliated stochastic Lie systems [6, 27]. Moreover, we also plan to study a stochastic Hamiltonian system appearing in celestial mechanics concerning the stochastic variation of the inertia tensor [64], biological methods, epidemiological models, coronavirus systems, etc. Additionally, it would be interesting to study other types of superposition rules appearing in stochastic differential equations [45] and to apply the Poisson coalgebra method to particular Hamiltonian stochastic Lie systems. Finally, we would like to analyse the potential extensions of methods designed for Lie systems, or their generalisations, to stochastic counterparts. For instance, PDE Lie systems appear in hydrodynamic equations in [39] and the analysis of conditional symmetries in [37]. It would be interesting to extend, first, the notion of a stochastic Lie system to PDE Lie systems [23, 63]. Moreover, one of the authors of this work along with his colleagues are studying hydrodynamical equations with a stochastic character, e.g. stochastic Burgers equations [39, 43], so as to generalise the Riemann invariant method [38] to a stochastic realm, and conditional and standard Lie symmetries for stochastic partial differential equations, partially attempting to extend the methods of [37] and, potentially, other works [59].